OpenStudy (izzy18970):

one last question for the day. please help!?!? write and solve an equation to solve each problem 3 consecutive even integers can be represented by n,n + 2, and n+ 4. If the sum of 3 consecutive integers is 36, what are the integers?

OpenStudy (izzy18970):

medal given for helping

OpenStudy (izzy18970):

I need help!?!?!?! please???

OpenStudy (anonymous):

n = 10

OpenStudy (izzy18970):

can you do that in steps so i can understand?

OpenStudy (anonymous):

n+(n+2)+(n+4) = 36 2n+2+n+4=36 3n+6=36 n=(36-6)/3 n=10

OpenStudy (anonymous):

give me medal lol

OpenStudy (izzy18970):

thank you so much 

OpenStudy (ranga):

They want the integers. So don't forget to give the answer as: 10, 12 and 14

OpenStudy (izzy18970):

?